# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

### v3

@@ 3.1 Relative frequencies
In the following, we analyze several statistics of multiperiodic sequences and multiperiodic processes. We begin with the relative frequencies and the marginal distribution.
@@ 3.2 Waiting times
Another important statistic of a stochastic process is the waiting time. We define the waiting time as
@@
The Kac lemma [45] states that
@@
for a stationary ergodic process $(K_{t})_{t\in\mathbb{Z}}$. However, for a multiperiodic process, we can obtain a stronger uniform bound.
@@
By the way, we notice this fact.
@@ 3.3 Number of types
We denote its cardinality
$V_{t}:=\operatorname{card}\mathcal{V}_{t}.$ (41)
Let us introduce also four more statistics:
@@
The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{card}\mathcal{V}_{t}$. ∎
Generalizing an idea used by Khmaladze [48] for memoryless sources, we can prove this proposition.
@@
Quotient $V_{t}/t$ is called the type-token ratio in quantitative linguistics [41, 48, 1]. It is easy to see that it tends to zero in the stationary case.
@@
We need a bit stronger bound.
@@
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [5, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
@@ 4.2 Block entropy
In this section, we sandwich bound the Shannon entropy in terms of the number of types. For this goal, we use a bound stemming from the Hölder and Jensen inequalities. Let $\ln$ denote the natural logarithm.
@@
$u_{t}-1\leq H(K_{1}^{t})\leq\left[e\operatorname{\mathbf{E}}V_{t}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2.$ (63)
@@
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left[q\log\operatorname{\mathbf{E}}M_{t}^{1/q}+\log t+\frac{\log e}{p-1}\right]+2$
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left[(q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right]+2.$ (66)
@@
$H(K_{1}^{t})$ $\leq\left[(t^{p-1}\operatorname{\mathbf{E}}V_{t})^{1/p}+3\right]\left[(q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right]+2$
$=\left[(e\operatorname{\mathbf{E}}V_{t})^{\frac{\ln t}{\ln t+1}}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2$
$\leq\left[e\operatorname{\mathbf{E}}V_{t}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2,$ (67)
@@ 5.1 Constant periods
In the following, we work out some concrete examples of processes. First, we analyze the statistics of the multiperiodic process with all periods equal to a fixed constant. This case leads to the number of types that grows logarithmically with the sample size.
Let $\pi_{k}=p$ for all $k\in\mathbb{N}$. Then we obtain
@@
$p_{k}$ $=f_{k}-f_{k+1}=\frac{1}{p}\left(\frac{p-1}{p}\right)^{k-1},$ (70)
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=2p\left(\frac{p}{p-1}\right)^{k-1}-1.$ (71)
@@
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}=\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+2,$ (72)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}=\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1,$ (73)
$\log q_{t}$ $=\log f_{g_{t}+1}=-g_{t}\log\left(\frac{p}{p-1}\right)\leq-\log(t+1)+\log(p).$ (74)
@@ 5.2 Linear periods
Second, we inspect the statistics of the multiperiodic process with periods that grow linearly. This case leads to the number of types that grows like a power law. The power-law growth of the number of types is known as the Herdan-Heaps law in quantitative linguistics [41, 39, 1].
Let $\pi_{k}=1+\left\lceil ck\right\rceil$ for a $c>0$ and all $k\in\mathbb{N}$. We may upper bound
@@
In consequence, we obtain these bounds
@@
$p_{k}$ $=f_{k}-f_{k+1}\leq\frac{1}{ck+1}\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\frac{1}{c}\left(\frac{c+2}{c}\right)^{1/c}k^{-(c+1)/c},$ (79)
@@
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\leq\left(\frac{t}{c}\left(\frac{c+2}{c}\right)^{1/c}\right)^{c/(c+1)}.$ (82)
We also need the converse bounds
$p_{k}$ $\geq\frac{1}{ck+2}(ek)^{-1/c}\geq\frac{e^{-1/c}}{c+2}k^{-(c+1)/c},$ (83)
$g_{t}$ $\leq\left(\frac{te^{-1/c}}{c+2}\right)^{c/(c+1)}$ (84)
@@
As a result, bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ implies the Herdan-Heaps law
@@ 6 Conclusion
The toy example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of redundancy for certain stationary sources, possibly entailing statistical language models. This topic is relevant not only for quantitative linguistics but also for large language models in machine learning, via the neural scaling law. As it was shown in the previous research revolving around the example of Santa Fe processes, the topic can be linked to more traditional analyses of Zipf‘s law. This study has supplied further findings in this vein.

### v4

@@ 3 Statistics
In this section, we analyze several statistics of multiperiodic sequences and multiperiodic processes. We begin with the relative frequencies and the marginal distribution. The waiting times are the second kind of statistics. The third sort of investigated quantities relate to the number of observed types. Some results concerning these quantities are established by applying the decimation properties shown in Theorem 2, the Birkhoff ergodic theorem, the union bound, and the Borel-Cantelli lemma.
@@ 3.1 Relative frequencies
By the decimation property established in Theorem 2, for a multiperiodic sequence there exist limiting relative frequences.
@@
Hence by ergodicity of the multiperiodic process, the limiting relative frequencies are equal to the marginal distribution of the process.
@@ 3.2 Waiting times
Another important statistic of a stochastic process is the waiting time. Following [65, 66], we define the waiting time as the first position in the process where the value of interest occurs. That is we put
@@
The celebrated Kac lemma [65] predicts that the conditional expectation of the waiting time equals the inverse probability of the recurring value. Namely, we have
@@
for a stationary ergodic process $(K_{t})_{t\in\mathbb{Z}}$.
Complementing this result, for a multiperiodic process, we can obtain a uniform bound as follows.
@@
Moreover, for increasing periods, we obtain that the waiting times increase monotonically, whereas the marginal probabilities are monotonically decreasing.
@@ 3.3 Number of types
Now we proceed to bounding the number of types observed in a finite sample. We prove results of a varying generality. Some of these propositions hold for general stationary sources and may be useful for quantitative linguistic applications revolving around studying the type-token ratio [50].
@@
We denote its cardinality, the random number of types in a finite sample,
$V_{t}:=\operatorname{\#}\mathcal{V}_{t}.$ (41)
To bound the number of types effectively, we introduce four other statistics:
@@
The first result is a simple sandwich bound for multiperiodic processes.
@@
The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{\#}\mathcal{V}_{t}$. ∎
The subsequent two statements pertain to general stationary processes. We note that functions $g_{t}$ and $q_{t}$ were introduced by Khmaladze [52] for memoryless sources. Generalizing his technique from memoryless processes to stationary ones, we can prove this proposition.
@@
Quotient $V_{t}/t$ is called the type-token ratio in quantitative linguistics [49, 50]. It is easy to see that it tends to zero in the stationary case.
@@
We will need a stronger bound that holds under a mild condition.
@@
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [61, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
@@ 4 Information
In this section, we demonstrate propositions that are central from our particular perspective. The first result concerns the flow of information in the description of a sample drawn from a multiperiodic process. We show that a sample of a multiperiodic process carries exactly the same information as the set of seeds for types appearing in the sample. These two objects can be computed from one another and hence their entropies are equal. Consequently, this result allows to sandwich bound the block entropy of a multiperiodic process by the number of types and related statistics. In particular, Hilberg‘s law follows from Heaps‘ law, cf. [4] for a similar result for Santa Fe processes.
@@ 4.2 Block entropy
Now, we will sandwich bound the Shannon entropy in terms of the number of types. For this goal, we will use a bound stemming from the Hölder and Jensen inequalities. Let $\ln$ denote the natural logarithm.
@@
$u_{t}-1\leq H(K_{1}^{t})\leq\left(e\operatorname{\mathbf{E}}V_{t}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2.$ (63)
@@
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left(q\log\operatorname{\mathbf{E}}M_{t}^{1/q}+\log t+\frac{\log e}{p-1}\right)+2$
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left((q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right)+2.$ (66)
@@
$H(K_{1}^{t})$ $\leq\left((t^{p-1}\operatorname{\mathbf{E}}V_{t})^{1/p}+3\right)\left((q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right)+2$
$=\left((e\operatorname{\mathbf{E}}V_{t})^{\frac{\ln t}{\ln t+1}}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2$
$\leq\left(e\operatorname{\mathbf{E}}V_{t}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2,$ (67)
@@ 5 Examples
Theorem 14 implies that Hilberg‘s law (4) for a multiperiodic process essentially follows from Heaps‘ law (9). In this section, we will work out two concrete examples of processes. First, we will analyze the multiperiodic process with all periods equal to a fixed constant. This case leads to the number of types that grows logarithmically with the sample size. Second, we will inspect the multiperiodic process with periods that grow linearly. This case leads to the number of types that grows like a power law, Heaps‘ law (9) indeed. Consequently, for this example, we also obtain Hilberg‘s law (4).
@@ 5.1 Constant periods
First, let us consider the multiperiodic process where periods equal to a fixed constant. To be concrete, let $\pi_{k}=p$ for all $k\in\mathbb{N}$. Then we obtain a geometric distribution of $P(K_{0}=k)=p_{k}$. In particular, we may bound
@@
$p_{k}$ $=f_{k}-f_{k+1}=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}=\frac{1}{p}\left(\frac{p-1}{p}\right)^{k-1},$ (70)
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=\sum_{j=1}^{k-1}\left(\frac{p}{p-1}\right)^{j}+p\left(\frac{p-1}{p}\right)^{k-1}$
$=\frac{\left(\frac{p}{p-1}\right)^{k}-1}{\frac{p}{p-1}-1}-1+\left(\frac{p}{p-1}\right)^{k-1}p=2p\left(\frac{p}{p-1}\right)^{k-1}-1.$ (71)
@@
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}=\min\left\{k:2p\left(\frac{p}{p-1}\right)^{k-1}>t+1\right\}$
$=\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+2,$ (72)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}=\max\left\{k:p\left(\frac{p}{p-1}\right)^{k-1}\leq t\right\}$
$=\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1,$ (73)
$\log q_{t}$ $=\log f_{g_{t}+1}=-g_{t}\log\left(\frac{p}{p-1}\right)\leq-\log(t+1)+\log p.$ (74)
@@ 5.2 Linear periods
Now, we consider the multiperiodic process where periods grow approximately in a linear fashion. To be concrete, let $\pi_{k}=1+\left\lceil ck\right\rceil$ for a $c>0$ and all $k\in\mathbb{N}$. We may upper bound
@@
In consequence, we obtain Zipf‘s law (7) for $P(K_{0}=k)=p_{k}$. In particular, we may bound
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\geq(ek)^{-1/c},$ (78)
@@
$p_{k}$ $=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\geq\frac{1}{ck+1}(ek)^{-1/c}\geq\frac{e^{-1/c}}{c+2}k^{-(c+1)/c},$ (80)
$p_{k}$ $=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\leq\frac{1}{ck+1}\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\frac{1}{c}\left(\frac{c+2}{c}\right)^{1/c}k^{-(c+1)/c}.$ (81)
Thus we may bound the waiting times
@@
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\leq\left(\frac{t}{c}\left(\frac{c+2}{c}\right)^{1/c}\right)^{c/(c+1)},$ (84)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\geq\left(\frac{te^{-1/c}}{c+2}\right)^{c/(c+1)}$ (85)
@@
As a result, bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ implies Heaps‘ law
@@ 6 Conclusion
The example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of entropy for certain stationary sources, possibly entailing statistical language models [3, 11]. This topic is relevant not only for quantitative linguistics [2, 34] but also for large language models in machine learning, via the neural scaling law [5, 6, 7, 8, 4]. As it was shown in the previous research revolving around the example of Santa Fe processes [2, 34, 35, 43], the topic can be linked to more traditional analyses of Zipf‘s law [36, 37]. This study has supplied further examples and findings. We hope that the earlier constructed Santa Fe processes and the multiperiodic processes described here may be applied as toy data sources to be compared with natural data in the study of language and artificial intelligence.